# Causal Inference with Missing Exposures and Missing Outcomes

### v2

@@ Introduction
Then given the observed data $O=(L_{0},\Delta_{A},A,\Delta_{Y_{0}},Y_{0},L_{1},\Delta_{Y_{1}},Y_{1})$, we can identify denominator and numerator, in turn. The denominator $\mathbb{P}(Y_{0}^{*}=0)$ represents the counterfactual prevalence of not having the outcome at baseline and, thus, being at risk. The causal structure for this parameter is analogous to that of Section 3.2, but with an additional measurement indicator for the baseline outcome. Therefore, under analogous assumptions, we can identify $\mathbb{E}(Y_{0}^{*})=\mathbb{E}\big[\mathbb{E}(Y_{0}\mid\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0})\big]$ with proof in Appendix S2.3. Since we are interested the counterfactual probability of being at risk at baseline $\mathbb{P}(Y_{0}^{*}=0)$, our statistical estimand for the denominator becomes $1-\mathbb{E}\big[\mathbb{E}(Y_{0}\mid\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0})\big]$.
In our final causal parameter (Eq. 1), the numerator $\mathbb{P}(Y_{1}^{*}=1,Y_{0}^{*}=0)$ represents the counterfactual probability of having the outcome at endline but not at baseline. For ease of notation, let $Z^{*}=\mathbb{I}(Y_{1}^{*}=1,Y_{0}^{*}=0)$ represent the joint indicator of these two counterfactual values. To identify $\mathbb{E}(Z^{*})=\mathbb{P}(Z^{*}=1)$, we need analogous assumptions as for the denominator together with the following. Among those known to be at risk at baseline ($\Delta_{Y_{0}}=1,Y_{0}=0$) and with measured exposure of interest ($\Delta_{A}=1,A=a$): the baseline and time-varying covariates capture the common causes of the joint outcome and endline measurement: $Z^{*}\perp\Delta_{Y_{1}}\mid L_{1},Y_{0}=0,\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0}$. We also need there to be a positive probability of follow-up measurement within all possible values of $L_{0}$ and $L_{1}$: $\mathbb{P}(\Delta_{Y_{1}}=1\mid L_{1},Y_{0}=0,\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0})>0$ a.e. Under these assumptions and with proof given in Appendix S2.3, the numerator is identified as
@@
Putting it all together, the statistical estimand with missing exposures, missing outcomes at baseline, and missing outcomes at endline is given by
@@
In the previous section, we introduced a series of causal models and identification results of increasing complexity. For the resulting statistical estimands, we could use a singly robust estimation approach, such as standardization (a.k.a., “G-computation”) or inverse probability weighting (IPW) [9, 13]. Here, we highlight the use of TMLE, which is a doubly robust estimation procedure and asymptotically efficient under certain conditions [11]. In TMLE, initial estimates of the relevant pieces of the observed data distribution are updated to achieve the optimal bias-variance trade-off for the estimand and to solve the efficient influence curve equation. Initial estimates are often computed via Super Learner, an ensemble machine learning algorithm using V-fold cross-validation to select an optimal weighted linear combination of predictions from a library of candidate learners [48]. Thereby, TMLE leverages machine learning to avoid introducing new modeling assumptions during estimation, while supporting valid statistical inference under reasonable conditions. Notably, for ratio-type estimands corresponding to Counterfactual Strata Effects (Eq. 2), we would implement a separate TMLE for the estimand in the numerator (the joint probability) and the estimand in the denominator (the marginal baseline probability) before combining the results.
TMLE is an asymptotically linear estimator and is normally distributed in the large data limit [11]. The estimator minus the estimand behaves like a sample mean in the first order: $\hat{\Psi}-\Psi=\frac{1}{N}\sum_{i=1}^{N}D_{i}+o_{P}(N^{-1/2})$ where $D_{i}$ is the influence curve for participant $i=\{1,\ldots,N\}$ and $o_{P}(N^{-1/2})$ is a second-order remainder term going to zero in probability [49]. The estimated influence curve is used to calculate standard errors, Wald-type confidence intervals, and p-values. Concretely, a 95% confidence interval is constructed using $\hat{\Psi}\pm z_{0.0975}\frac{\hat{\sigma}}{\sqrt{n}}$ where $z_{0.975}$ is the critical value at the 97.5th-percentile of the standard normal and $\hat{\sigma}$ is the standard deviation of the estimated influence curve. For ratio-type estimands (Eq. 2), once the influence curves for the numerator and denominator have been estimated, the Delta method provides an estimate of the influence curve for our overall estimand. Then to calculate measures of association on the difference, ratio, or odds ratio scale, we apply the Delta method a second time to get inference for these types of functionals.
@@
We now return to our motivating question: what is the effect of alcohol use on incident TB infection among adults in rural Eastern Uganda? With our multinational and interdisciplinary team, we worked through the Causal Roadmap to specify the Statistical Analysis Plan, including the causal model and the adjustment sets [21, 22, 23, 50, 38]. Concretely, the causal model reflected team’s knowledge of study protocol, the study procedures, and the epidemiology of TB in the region. Our adjustment set included the SEARCH trial arm, community indicators, household HIV status, as well as individual-level age, sex, and mobility measures. For the primary analysis, we used TMLE with Super Learner to combine estimates from generalized linear models, multivariate adaptive regression splines, and the mean. We conducted influence curve-based inference, accounting for clustering by household (Appendix S4) [51, 26]. In secondary analyses, we considered communities, instead of households, to be the independent unit. To examine the sensitivity of our results to alternative estimation approaches, we also implemented IPW for the same statistical estimand, but using parametric regressions to estimate the weights. Finally, to examine the impact of handling of our missing data, we took the following “naïve” approach: (1) subset on participants known to be at risk of TB at baseline ($Y_{0}=0$); (2) further subset on participants with measured exposures ($\Delta_{A}=1$) and measured outcomes at endline ($\Delta_{Y_{1}}=1$), and (3) implement TMLE to adjust for confounding and to estimate $\mathbb{E}[\mathbb{E}(Y_{1}\mid A=a,L)\mid Y_{0}=0]$. This approach is inherently flawed, because of the bias induced by inappropriately conditioning on various colliders and mediators (Figure 3). Nonetheless, the approach could be one taken by an analyst aiming to implement a “complete-case” analysis while adjusting for confounding [52, 53, 54]. For statistical inference, both IPW and the naïve approach accounted for clustering by household.
@@
In the primary analysis using TMLE with clustering by household, we found that alcohol use was associated with a 49% increase in the risk of incident TB: risk ratio (RR)=1.49 (95%CI: 1.39-1.59) [38]. As shown in Figure 4, secondary analyses with the community as the independent unit yielded very similar results, despite meaningfully reducing the effective sample size from 1,380 households to 9 communities: RR=1.49 (95%CI: 1.37-1.62); Appendix S4. In contrast, IPW relying on parametric assumptions resulted in a smaller association and confidence intervals overlapping the null: RR=1.13 (95%CI: 1.00-1.27). Finally, after restricting to participants who tested negative at baseline, responded to questions about alcohol use, and tested again at follow-up, the naïve approach was the least precise and resulted in the widest confidence intervals: RR=1.18 (95%CI: 0.89-1.57).
@@
Figure 4: Results from SEARCH-TB for the association of alcohol use on incident tuberculosis (TB) infection: “Primary” with TMLE and clustering by household, “Secondary” with TMLE and clustering by community, “IPW” with inverse probability weighting, and “Naïve” based on subsetting on those at risk at baseline and with measured exposures and outcomes at endline.
@@ Discussion
We presented causal models, causal parameters, and identification results for a series of prospective studies with increasing levels of missingness. For estimation and inference, we highlighted the use of TMLE with Super Learner to robustly and efficiently estimate the corresponding statistical estimands. Application to real-data from SEARCH-TB demonstrated the real-world consequences of our work. Using TMLE to flexibly account for confounding and missingness, we found a 49% relative increase in the risk of incident TB infection associated with drinking alcohol: RR=1.49 (95%CI: 1.39-1.59). For the same statistical estimand but using parametric regressions, IPW resulted in a smaller association and meaningfully wider confidence intervals: RR=1.13 (95%CI: 1.00-1.27). Finally, the naïve approach to condition on the participants with baseline TB risk and measured exposures/outcomes yielded a null association (RR=1.18, 95%CI: 0.89-1.57).
@@
There are several advantages to our framework for handling missing exposures and outcomes when evaluating the causal effects. Unlike the m-DAG approach which aids in identifying the entire joint distribution of the observed data [4], our approach focuses our efforts on the portions of the observed data distribution relevant for the statistical estimand. Furthermore, our approach uses data on all participants, improving efficiency relative to approaches excluding participants with missing data on the relevant variables (sometimes called an “available-case analysis”) [4]. Finally, our approach leads us to statistical estimands that can be robustly and rigorously estimated with modern methods, such as TMLE.
A major limitation to this work is that we did not address missingness on the confounders. Some approaches include mean/median imputation or including missingness indicators in the adjustment set. For further discussion on missing confounders, we refer to the comprehensive work of Williamson et al. [55]. We also focused on cross-sectional or prospective studies. Thus, we did not cover scenarios where the outcome impacts the measurement of other variables. Such scenarios would arise in case-control studies and have been addressed in prior literature (e.g., [56, 57, 58, 59, 54]).
@@
There are other limitations to our work. First, we did not provide an exhaustive set of causal models and identification results for all possible studies; however, our approach is generalizable and covers many scenarios arising in public health. Second, in our real-data demonstration, we did not include multiple imputation, which is a common approach for missing data and can also leverage machine learning [60, 61, 62]. Future work is needed to investigate the assumptions, implementation, and performance of multiple imputation in settings mirroring our motivating example: (1) missingness on the exposure of interest, (2) missingness on the baseline outcome, (3) missingness in the final outcome, (4) confounding, and (5) dependence among study participants. Finally, we relied on various versions of the sequential exchangeability assumption for both confounding and missingness. In practice, data may be missing as a result of unobserved variables, and we may need to collect additional data as well as conduct sensitivity/bias analyses [63, 64]. Nonetheless, even when a “causal gap” remains, we have a framework to define a statistical estimand, which is aligned with our research question [21, 65, 23, 50].
Overall, our goal was to demonstrate how an established framework for estimating causal effects with censored outcomes could be extended to settings with missingness on the exposure, the outcome at baseline, and the outcome at endline. To do so, we provided a series of causal models with increasing complexity and discussed the identification assumptions for each. For estimation and inference, we focused on TMLE, a doubly-robust approach using machine learning and data from all participants (including those with missing exposures or outcomes). Our work was motivated by the investigation of the effect of alcohol use on incident TB infection [38], and our re-analysis demonstrated the real-world consequences of our approach.
Source of Funding: This work was supported, in part, by The National Institutes of Health (awards: R01AI151209 (CM), K23AI118592 (CM), U01AI099959, and UM1AI068636), the President’s Emergency Plan for AIDS, and the AIDS Research Institute at the University of California San Francisco.
Conflicts of Interest: The authors report no conflict of interest in this work.
Acknowledgments: We thank the Ministries of Health of Uganda and Kenya; our research and administrative teams in San Francisco, Uganda, and Kenya; our collaborators and advisory boards; and, especially, all the communities and participants involved. We also thank Dr. Diane Havlir and Dr. Maya L. Petersen, who together with Dr. Moses R. Kamya are the MPIs of the SEARCH collaboration.
Data and Code Availability: A de-identified dataset and computing code sufficient to reproduce the study findings will be made available following approval of a concept sheet summarizing the analyses to be done. Further inquiries can be directed to the SEARCH Scientific Committee at douglas.black@ucsf.edu.
@@
  • Appendix S1: More on the classic point-treatment problem
@@
Appendix S1: More on the classic point-treatment problem
@@
Under interventions on the causal model, we generate counterfactual outcomes corresponding to the research question of interest. Specifically, let $Y^{a}$ be the counterfactual outcome for a given participant if, possibly contrary-to-fact, they had exposure-level $A=a$. Then our causal target parameter $\mathbb{E}(Y^{a})$ is the counterfactual mean outcome if all study participants had exposure-level $A=a$. In our running example, $\mathbb{E}(Y^{a})=\mathbb{P}(Y^{a}=1)$ is the counterfactual risk of incident TB infection with alcohol use $A=a$.
@@
To identify our causal target parameter and express it as function of the distribution of the observed data $O=(L,A,Y)$, we would need there to be no unmeasured confounding, which corresponds to the assumption that the baseline covariates $L$ capture all the joint causes of the exposure $A$ and outcome $Y$. This condition is called “the randomization assumption” or “exchangeability”, can be evaluated with the backdoor criterion, and can be represented as $Y^{a}\perp A\mid L$ [9, 10, 11, 12].
Concretely, $\mathbb{E}(Y^{a})$ is not identified in Figure 5A because there is an unmeasured common cause, represented by $U^{*}$, of the exposure $A$ and outcome $Y$. In Figure 5B and C, we show two causal models where this assumption would hold.
Additionally, we need there to be a non-zero probability of having the exposure in all possible values of $L$: $\mathbb{P}(A=a\mid L)>0$ a.e. This is a condition on data support and known as the “positivity assumption”. We note that the consistency assumption, stating that the counterfactual outcome $Y^{a}$ equals the observed outcome $Y$ under exposure $A=a$, holds by design when we define counterfactuals through intervention on the causal model. Under the exchangeability and positivity assumptions, our causal target is equal to the statistical estimand given by the G-computation formula: $\mathbb{E}[\mathbb{E}(Y\mid A=a,L)]$ [9]. Even if these assumptions are not reasonable (e.g., Figure 5A reflects reality), we still have well-defined statistical estimand, on which we can focus our estimation efforts. In other words, we still proceed with estimation and inference $\mathbb{E}[\mathbb{E}(Y\mid A=a,L)]$ , while appropriately accounting for lack of identification in our interpretation [21, 22, 23, 66, 65] For ease of presentation, we provide the causal models without dependence between the unmeasured factors in the main text.
@@
$=$ $\sum_{l_{0}}\mathbb{P}(Y^{*}=1\mid\Delta_{A}=1,A=a,L_{0}=l_{0})\mathbb{P}(L_{0}=l_{0})$
@@
$=$ $\sum_{l_{0}}\sum_{l_{1}}\mathbb{P}(Y^{*}=1\mid\Delta_{Y}=1,L_{1}=l_{1},\Delta_{A}=1,A=a,L_{0}=l_{0})\times$
$\quad\mathbb{P}(L_{1}=l_{1}\mid\Delta_{a}=1,A=a,L_{0}=l_{0})\mathbb{P}(L_{0}=l_{0})$
@@
Let $Y_{0}^{*}=Y_{0}^{\Delta_{A}=1,A=a,\Delta_{Y_{0}}=1}$ and $Y_{1}^{*}=Y_{1}^{\Delta_{A}=1,A=a,\Delta_{Y_{0}}=1,\Delta_{Y_{1}}=1}$. Recall that we defined the target parameter for this section as

### v3

@@ Introduction
Then given the observed data $O=(L_{0},\Delta_{A},A,\Delta_{Y_{0}},Y_{0},L_{1},\Delta_{Y_{1}},Y_{1})$, we can identify the denominator and the numerator. The denominator $\mathbb{P}(Y_{0}^{*}=0)$ represents the counterfactual prevalence of not having TB at baseline — under alcohol use $A=a$ and under complete measurement of the exposure and baseline outcome ($\Delta_{A}=\Delta_{Y_{0}}=1)$. The causal structure for this parameter is analogous to that of Section 3.2, but with an additional measurement indicator for the baseline outcome. Therefore, under analogous assumptions, we can identify $\mathbb{E}(Y_{0}^{*})=\mathbb{E}\big[\mathbb{E}(Y_{0}\mid\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0})\big]$ with proof in Appendix S3.3. Since we are interested the counterfactual probability of being at-risk at baseline $\mathbb{P}(Y_{0}^{*}=0)$, our statistical estimand for the denominator becomes $1-\mathbb{E}\big[\mathbb{E}(Y_{0}\mid\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0})\big]$. Altogether, this statistical estimand accounts for differential measurement of the exposure and baseline TB status as well as the impact of alcohol use $A=a$ on baseline TB risk.
In our final causal parameter (Eq. 1), the numerator $\mathbb{P}(Y_{1}^{*}=1,Y_{0}^{*}=0)$ represents the counterfactual probability of acquiring TB — under alcohol use $A=a$ and under complete measurement $\Delta_{A}=\Delta_{Y_{0}}=\Delta_{Y_{1}}=1$. For ease of notation, let $Z^{*}=\mathbb{I}(Y_{1}^{*}=1,Y_{0}^{*}=0)$ represent the joint indicator of these two counterfactual values. To identify $\mathbb{E}(Z^{*})=\mathbb{P}(Z^{*}=1)$, we need analogous assumptions as for the denominator together with the following. Among those known to be at-risk at baseline ($\Delta_{Y_{0}}=1,Y_{0}=0$) and with measured exposure of interest ($\Delta_{A}=1,A=a$): the baseline and time-varying covariates capture the common causes of the joint outcome and endline measurement: $Z^{*}\perp\Delta_{Y_{1}}\mid L_{1},Y_{0}=0,\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0}$. We also need there to be a positive probability of follow-up measurement within all possible values of $L_{0}$ and $L_{1}$: $\mathbb{P}(\Delta_{Y_{1}}=1\mid L_{1},Y_{0}=0,\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0})>0$ a.e. Under these assumptions and with proof given in Appendix S3.3, the numerator is identified as
@@
Putting it all together, our statistical estimand is given by
@@
In the previous section, we introduced a series of causal models and identification results of increasing complexity. For the resulting statistical estimands, we could use singly robust estimators, such as standardization (a.k.a., “G-computation”) or inverse probability weighting (IPW).[61, 31] Here, we highlight the use of TMLE, which is a doubly robust procedure and asymptotically efficient under certain conditions.[76] In TMLE, initial estimates of the relevant pieces of the observed data distribution are updated to achieve the optimal bias-variance trade-off for the estimand and to solve the efficient influence curve equation. Initial estimates are often computed via Super Learner, an ensemble machine learning algorithm using V-fold cross-validation to select an optimal weighted linear combination of predictions from a library of candidate learners.[75] Thereby, TMLE leverages machine learning to avoid introducing new modeling assumptions during estimation, while supporting valid statistical inference under reasonable conditions. Notably, for ratio-type estimands corresponding to Counterfactual Strata Effects (Eq. Appendix S4.1: Simulation Results), we would implement a separate TMLE for the numerator estimand (the joint probability) and the denominator estimand (1- the baseline prevalence) before combining the results.
TMLE is an asymptotically linear estimator and is normally distributed in the large data limit.[76] The estimator minus the estimand behaves like a sample mean in the first order: $\hat{\Psi}-\Psi=\frac{1}{N}\sum_{i=1}^{N}D_{i}+o_{P}(N^{-1/2})$ where $D_{i}$ is the influence curve for participant $i=\{1,\ldots,N\}$ and $o_{P}(N^{-1/2})$ is a second-order remainder term going to zero in probability.[79] The estimated influence curve is used to calculate standard errors, Wald-type confidence intervals, and p-values. Concretely, a 95% confidence interval is constructed using $\hat{\Psi}\pm z_{0.0975}\frac{\hat{\sigma}}{\sqrt{n}}$ where $z_{0.975}$ is the critical value at the 97.5th-percentile of the standard normal and $\hat{\sigma}$ is the standard deviation of the estimated influence curve. For ratio-type estimands (Eq. Appendix S4.1: Simulation Results), once the influence curves for the numerator and denominator have been estimated, the Delta method provides an estimate of the influence curve for our overall estimand.[79, 45] Then to calculate measures of association on the difference, ratio, or odds ratio scale, we apply the Delta method a second time to get inference for these types of functionals (Appendix S4).
To evaluate the practical performance of our approach, we conducted a simulation study reflecting the motivating study (Figure 3). The simulation design and results are detailed in Appendix S5. Briefly, we found that TMLE had negligible bias and achieved nominal confidence interval coverage for the exposure-specific estimands $\Psi(\mathbb{P};a)$, given in Eq. Appendix S4.1: Simulation Results for $A=\{0,1\}$, as well as their ratio (i.e. the relative effect) and their difference (i.e., the average treatment effect). All other approaches exhibited meaningful bias and sub-nominal confidence interval coverage.
@@
We now return to our motivating question: what is the effect of alcohol use on incident TB infection among adults in rural Eastern Uganda? With our multinational and interdisciplinary team, we worked through the Causal Roadmap to specify the Statistical Analysis Plan.[53, 15, 14, 21, 48, 1] Our causal model reflected the team’s knowledge of the study protocol, study procedures, and epidemiology of TB in the region. Our adjustment set included the SEARCH trial arm, community indicators, household HIV status, as well as individual-level age, sex, and mobility measures. The exposure was self-report of any alcohol use, and the outcome was incident TB infection: conversion from a negative tuberculin skin test to a positive one.[1] The covariates and exposure were measured at the start of the parent SEARCH study (2013-2014), while TB status was measured subsequently in the sub-study (2015-2017).[25, 39, 40] For the primary analysis, we used TMLE with Super Learner to combine estimates from generalized linear models, multivariate adaptive regression splines, and the mean. We conducted influence curve-based inference, accounting for clustering by household (Appendix S6).[78, 50]
In secondary analyses, we considered communities, instead of households, to be the independent unit. To examine the sensitivity of our results to alternative estimation approaches, we implemented two singly robust approaches for the final statistical estimand (Eq. Appendix S4.1: Simulation Results): G-Computation using parametric regressions to estimate the iterated conditional expectations and IPW using parametric regressions to estimate the weights.[8] Finally, to examine the impact of our handling of missing data, we took the following “naïve” approach: (1) subset on participants with complete data $\Delta_{A}=\Delta_{Y_{0}}=\Delta_{Y_{1}}=1$; (2) further subset on participants known to be at-risk at baseline $Y_{0}=0$, and (3) implement TMLE to adjust for confounding and to estimate $\mathbb{E}[\mathbb{E}(Y_{1}\mid A=a,L)]$ among the remaining subset of participants. This approach is inherently flawed, because of the bias induced by inappropriately conditioning on various colliders and mediators (Figure 3). Nonetheless, the approach could be one taken by an analyst aiming to implement a “complete-case” analysis while adjusting for confounding.[64, 41, 17] For statistical inference, G-computation, IPW, and the naïve approach accounted for clustering by household.
@@
In the primary analysis using TMLE with clustering by household, we found that alcohol use was associated with a 49% increase in the risk of incident TB: risk ratio (RR)=1.49 (95%CI: 1.39-1.59).[1] As shown in Figure 4, secondary analyses with the community as the independent unit yielded very similar results, despite meaningfully reducing the effective sample size to 9 communities: RR=1.49 (95%CI: 1.37-1.62); Appendix S6. In contrast, G-computation resulted in a larger association with large confidence intervals, while IPW resulted in a smaller association and confidence intervals overlapping the null: RR=1.58 (95%CI: 1.36-1.83) and RR=1.13 (95%CI: 1.00-1.27), respectively. Finally, after restricting to participants who responded to questions about alcohol use, tested negative at baseline, and tested again at follow-up, the naïve approach was the least precise and resulted in the widest confidence intervals: RR=1.18 (95%CI: 0.89-1.57).
@@
Figure 4: Results from SEARCH-TB for the association of alcohol use on incident tuberculosis (TB) infection: “Primary” with TMLE and clustering by household, “Secondary” with TMLE and clustering by community, “G-Comp” with G-computation, “IPW” with inverse probability weighting, and “Naïve” based on subsetting on those at-risk at baseline and with measured exposures and outcomes at endline.
@@ Discussion
We presented causal models, causal estimands, and identification results for a series of prospective studies with increasing levels of missingness. We reviewed how the Causal Roadmap could be applied to address missing outcomes and extended to address missing exposures. We then motivated the use of Counterfactual Strata Effects to precisely define causal estimands when the focus population is subject to missingness and/or impacted by the exposure. Specifically, we considered the common scenario where there is missingness on the outcome at baseline — determining who is at-risk of acquiring the outcome by endline. We further allowed for the exposure to influence baseline outcome risk. We demonstrated that while these real-world complications would result in “failures” in the target trial emulation framework,[26] our Roadmap approach facilitates definition and identification of the corresponding causal estimand. For estimation and inference, we highlighted the use of TMLE with Super Learner to robustly and efficiently estimate the corresponding statistical estimands.
We emphasize that Counterfactual Strata Effects are distinct from Principal Strata Effects.[16, 18, 19, 51] For simplicity, suppose that we did not have missingness, and define $Y_{0}^{a}$ and $Y_{1}^{a}$ as the counterfactual outcome at baseline and endline under exposure $A=a$. Then in SEARCH-TB, principal stratification could be applied to define the effect of alcohol on incident TB among the unobservable subset of participants who would have always been at-risk of TB at baseline regardless of their alcohol use: $\mathbb{P}(Y_{1}^{1}=1\mid Y_{0}^{1}=Y_{0}^{0}=0)$ versus $\mathbb{P}(Y_{1}^{0}=0\mid Y_{0}^{1}=Y_{0}^{0}=0)$. Instead, Counterfactual Strata Effects enable us to define the effect of alcohol on incident TB among the counterfactual population of persons at-risk at baseline under alcohol use $A=a$: $\mathbb{P}(Y_{1}^{1}=1\mid Y_{0}^{1}=0)$ versus $\mathbb{P}(Y_{1}^{0}=1\mid Y_{0}^{0}=0)$. The at-risk population is exposure-specific, and our estimand captures the effect of alcohol use on prevalent TB at baseline and on incident TB by endline. Of course, we could also directly examine the effect on prevalent TB at baseline: $\mathbb{P}(Y_{0}^{1}=1)$ versus $\mathbb{P}(Y_{0}^{0}=1)$. In addition to incidence studies,[40, 50] Counterfactual Strata Effects commonly arise when studying care cascades: the series of steps from screening, through diagnosis and treatment, to control.[7, 54, 2, 25, 5, 4, 47] Indeed, these cascades are a series of conditional probabilities — each subject to missingness and each potentially impacted by the exposure or intervention strategy.
@@
There are several advantages to our framework for handling multi-source missingness. As previously discussed, our approach avoids a common “failure” arising when aiming to emulate at target trial.[26] Furthermore, unlike the m-DAG approach which aids in identifying the entire joint distribution of the observed data,[46] our approach focuses our efforts on the portion of the observed data distribution relevant for the statistical estimand (Appendix S7). Our approach also uses data on all participants, improving efficiency relative to approaches excluding participants with missing data on the relevant variables (Figure 4; Appendix S.5). Finally, our approach leads us to statistical estimands that can be robustly and rigorously estimated with modern methods, such as TMLE.
@@
There are several limitations to our work. First, we focused on cross-sectional or prospective studies. Thus, we did not cover scenarios where the outcome impacts the measurement of other variables. Such scenarios would arise in case-control studies and have been addressed in prior literature.[65, 63, 86, 34, 17, 57] Second, we did not provide an exhaustive set of causal models and identification results for all possible studies; however, our approach is generalizable and covers many scenarios arising in public health. Third, we did not consider multiple imputation, which is a common approach for missing data.[66] However, foundational work is needed to investigate the assumptions, implementation, and performance of multiple imputation in settings mirroring our motivating example: (1) missingness on the exposure, (2) missingness on the baseline outcome, (3) missingness in the final outcome, (4) confounding, and (5) dependence among study participants. Indeed, Tompsett et al. suggested that multiple imputation cannot be used in settings with missingness on the focus population (termed “missing eligibility data”) and missingness on the exposure.[71] Finally, we relied on various versions of the sequential exchangeability assumption for both confounding and missingness. In practice, data may be missing as a result of unobserved variables, and we may need to collect additional data as well as conduct sensitivity/bias analyses.[36, 13] Nonetheless, even when a “causal gap” remains, we have a framework to define a statistical estimand, which is aligned with our research question, and appropriately addresses many real-world complications.[53, 15, 14, 21, 48]
@@
  • Appendix S1: The classic point-treatment problem
  • Appendix S2: Missing confounders
@@
  • Appendix S4: More on the Delta method
  • Appendix S5: Finite sample simulation study
@@
  • Appendix S7: Missingness DAGs
Appendix S1: The classic point-treatment problem
@@
Under interventions on the causal model, we generate counterfactual outcomes corresponding to the research question of interest. Specifically, let $Y^{a}$ be the counterfactual outcome for a given participant if, possibly contrary-to-fact, they had exposure-level $A=a$. Then our causal estimand $\mathbb{E}(Y^{a})$ is the expected counterfactual outcome if all had exposure-level $A=a$. In our running example, $\mathbb{E}(Y^{a})=\mathbb{P}(Y^{a}=1)$ is the counterfactual risk of incident TB infection with alcohol use $A=a$.
@@
To identify our causal target parameter and express it as function of the distribution of the observed data $O=(L,A,Y)$, we would need there to be no unmeasured confounding, which corresponds to the assumption that the baseline covariates $L$ capture all the joint causes of the exposure $A$ and outcome $Y$. This condition is called “the randomization assumption” or “exchangeability”, can be evaluated with the backdoor criterion, and can be represented as $Y^{a}\perp A\mid L$.[61, 52, 76, 28] Concretely, $\mathbb{E}(Y^{a})$ is not identified in Figure 5A because there is an unmeasured common cause, represented by $U^{*}$, of the exposure $A$ and outcome $Y$. In Figure 5B and C, we show two causal models where this assumption would hold.
Additionally, we need there to be a non-zero probability of having the exposure in all possible values of $L$: $\mathbb{P}(A=a\mid L)>0$ a.e. This is a condition on data support and known as the “positivity assumption”. We note that the consistency assumption, stating that the counterfactual outcome $Y^{a}$ equals the observed outcome $Y$ under exposure $A=a$, holds by design when we define counterfactuals through intervention on the causal model. Under the exchangeability and positivity assumptions, our causal target is equal to the statistical estimand given by the G-computation formula: $\mathbb{E}[\mathbb{E}(Y\mid A=a,L)]$.[61] Even if these assumptions are not reasonable (e.g., Figure 5A reflects reality), we still have well-defined statistical estimand, on which we can focus our estimation efforts. In other words, we still proceed with estimation and inference $\mathbb{E}[\mathbb{E}(Y\mid A=a,L)]$ , while appropriately accounting for lack of identification in our interpretation.[53, 15, 21, 82, 14] For ease of presentation, we provide the causal models without dependence between the unmeasured factors in the main text.
Appendix S2: Missing Confounders
Throughout the main text, we also assume that the covariates are completely measured. While missingness on covariates is common, a detailed discussion is beyond the scope of this manuscript, which focuses on defining, identifying, and estimating causal effects in settings with missing exposures and missing outcomes. Here, we provide a brief review of methods to handle missing confounders. Common approaches include dropping covariates subject to missingness, conducting a complete (available) case analysis, including a missingness indicator, and using multiple imputation.[20, 83, 81, 57] Excluding covariates subject to missingness will induce bias if those covariates are confounders. A complete case analysis, where observations with missing data are excluded, is guaranteed to reduce power and will be biased unless missingness is completely at random. In the missing indicator method, a dummy variable is included to indicate the covariate is measured/missing. Recently, the missingness indicator method was shown to perform similarly to multiple imputation (MI) in many settings.[83] MI uses observed variables to predict the values of missing ones and has traditionally relied on strong modeling assumptions, although there is a growing interest in using machine learning.[67, 35, 10]
Williamson et al. provide an excellent discussion and evaluation of more modern methods for missing confounders, including doubly robust approaches of generalized raking and targeted minimum loss-based estimation (TMLE).[81] These approaches are more robust to model misspecification and are efficient in certain scenarios. Williamson et al. also provide guidance on selecting the approach to handle missing confounders based on the bias-variance trade-off.[81] Recently, Qiu et al. extended this work and showed TMLE, a plug-in estimator, had superior performance to generalized ranking in finite samples.[57]
@@
$=$ $\sum_{l_{0}}\mathbb{P}(Y^{*}=1\mid A=a,\Delta_{A}=1,L_{0}=l_{0})\mathbb{P}(L_{0}=l_{0})$
@@
$=$ $\sum_{l_{0}}\sum_{l_{1}}\mathbb{P}(Y^{*}=1\mid\Delta_{Y}=1,L_{1}=l_{1},A=a,\Delta_{A}=1,L_{0}=l_{0})\times$
$\quad\mathbb{P}(L_{1}=l_{1}\mid A=a,\Delta_{A}=1,L_{0}=l_{0})\mathbb{P}(L_{0}=l_{0})$
@@
Let $Y_{0}^{*}=Y_{0}^{\Delta_{A}=1,A=a,\Delta_{Y_{0}}=1}$ and $Y_{1}^{*}=Y_{1}^{\Delta_{A}=1,A=a,\Delta_{Y_{0}}=1,\Delta_{Y_{1}}=1}$. Recall that we defined the target parameter for this section using Counterfactual Strata Effects:[3, 54, 5, 4, 50, 55, 22, 25, 40]